Consider the following data for two products of Gitano Manufacturing. (Loss amounts should be indicated with a minus sign. Ro1. Using direct labor hours as the basis for assigning overhead costs, determine the total production cost per unit for each1. Assign these three overhead cost pools to each of the two products using ABC. Overhead Cost Activity Drivers Activity Rate3. Consider the following additional information about these two product lines. If ABC is used for assigning overhead costs tProduct A Product B Total manufacturing costs Direct Materials per unit Direct Labor per unit Overhead per unit Total manufac
Answer #1
Products AProducts B





Number of units produced120001500Units


Direct Labor Cost @ 29 per DLH0.170.28DLH per unit


Direct Material Cost1.73Per Unit









ActivityOverhead Cost




Machine Setup67680




Material Handling54000




Quality Control107100





228780










1)Using Direct Labor Hours as basis for assigning Overhead Cost




ProductUnitsDLH per unitTotal Direct Labor Hours


A120000.172040


B15000.28420



13500
2460


Overhead Cost=$228780





Overhead cost per Direct Labor Hours=($228780/2460)=
93DLH








Overhead AssignedActivity DriverPredetermined Overhead RateTotal Overhead CostUnits ProducedOverhead Cost Per Unit

Product ADLH931897201200015.81

Product BDLH9339060150026.04








Total Manufacturing Cost of Product A





Direct Material Cost=12000*1.720400




Direct Labor Cost=2040*2959160




Overhead Cost=2040*$93189720




Total Manufacturing Cost269280




Manufacturing Cost per unit=$269280/1200022.44











Total Manufacturing Cost of Product B





Direct Material Cost=1500*34500




Direct Labor Cost=420*2912180




Overhead Cost=420*$9339060




Total Manufacturing Cost55740




Manufacturing Cost per unit=$55740/150037.16










2)If market price of product A is $28.68 and Product B is $56 Profit or loss



Product A





Price Per Unit28.68




Cost per Unit22.44




Profit or Loss6.24











Product B





Price Per Unit56




Cost per Unit37.16




Profit or Loss18.84










3)Products A
Product BTotal









Number of Setup required for Production32427


Number of Parts required18927


Inspection Hours required60255315









Machine Setup=67680/27 setup$            2,506.67Per Setup



Material Handling=54000/27 Parts$            2,000.00Per Unit



Quality Control=$107100/315$                340.00Inspection Hours










Product A





Direct Material Cost$          20,400.00




Direct Labor Cost$          59,160.00




Overhead





Setup=(3*2506.67)$            7,520.00




Parts=(18*2000)$          36,000.00




Quality Control=(60*340)$          20,400.00




Total Cost$       143,480.00




Cost Per Unit=$143480/12000$                  11.96











Product B





Direct Material Cost$            4,500.00




Direct Labor Cost$          12,180.00




Overhead





Setup=24*2506.67$          60,160.00




Parts=9*2000$          18,000.00




Quality Control=(255*340)$          86,700.00




Total Cost$       181,540.00




Cost Per Unit=$181540/1500$                121.03











Product A





Selling Price$                  28.68




Cost Per Unit$                  11.96




Profit or loss$                  16.72











Product B





Selling Price$                  56.00




Cost Per Unit$                121.03




Profit or loss$                (65.03)











Product A should be contine,while price of the product B should be increased.









4 (2)Yes





answered by: ANURANJAN SARSAM